# NeuCo-Bench: A Novel Benchmark Framework for Neural Embeddings in Earth Observation

### v1

@@ 5 Benchmark Evaluation
We describe how we tested NeuCo-Bench in a real‐world setting via CVPR EARTHVISION 2025, by detailing the experimental setup in Section 5.1, as well as discussion of outcomes and learnings in Section 5.2. We further present baseline evaluations exploring our set of downstream tasks in Section 5.3.
@@ 5.1 Data Challenge Validation
To validate NeuCo-Bench under realistic conditions, we utilized it in the 2025 CVPR EARTHVISION data challenge. Participants were tasked with compressing multi‐modal, multi‐temporal EO imagery, cf. Section˜4, into 1,024-dimensional embeddings. Given the SSL4EO-S12 input data cubes, this amounts for a compression ratio of approx. 7,000. Crucially, participants did not know which or the number of downstream tasks their embeddings would be evaluated on; this hidden‐task design discourages overfitting and encourages the development of general‐purpose EO representations. Participants were ranked, according to Section˜3, across two sets of downstream tasks. One modification was made to the dataset compared to the dataset described in Section˜4; the clouds task targets were mainly replaced by zeros, causing heavy skewness in the labels and basically random connection between the imagery and labels.
Phases. In a three‐week development phase, teams developed embedding methods using the publicly available SSL4EO v1.1 dataset. A partial release of 5 tasks, each with a subset of its samples but no information of the task type, allowed participants to receive initial feedback for development. Submissions returned only the mean $Q$ value to prevent leakage of task‐specific performance and information.
@@
Platform and infrastructure. The NeuCo-Bench framework was modified to utilize Eval.AI (Yadav et al., ) to collect submissions. NeuCo-Bench ran on a separate 8-vCPU server, retrieved new submissions via API, executed the evaluation, and pushed results to a custom leaderboard hosted on GitHub, displaying the dynamic ranking described in Section˜3, as well as back to Eval.AI. Additional details are available in Section˜A.2.
@@ 5.2 Data Challenge Results
Participation and ranking. Twenty‐three teams submitted to the development phase; sixteen went on to the final evaluation, nine of which shared their submissions publicly. The quality scores $Q_{t}^{(p)}$, shown in Fig.˜3 for the evaluation phase, varied widely from 0–5 on some tasks to 5–40 on others. The evaluation method efficiently scaled task importance ensuring that the tasks impacted the leaderboard relative to their differentiating effect across the teams. Notably, the weighting reduced the impact of the tasks with zero/random labels. Further, the dynamic ranking caused a swap between the original first and second place methods due to a third team, highlighting the impact of adaptive weighting.
Figure 3: Quality score $Q_{t}^{(p)}$ of the participants of the data challenge evaluation phase with corresponding task weight used for ranking. Teams are ordered by their final leaderboard rank, with the winner as Team 1. Team 10 is a simple averaging baseline described in Section˜5.3 and 11 denotes a baseline of normally distributed random embeddings.
Top methods. The team that achieved the best overall rank and the team with the highest average $Q$-Score both built their embeddings by ensembling multiple FM representations, although with different approaches; One pre-training backbones, and the other training a bottleneck based on frozen FM backbones. The fourth-place team took a different path, forgoing any pre-training and instead generating embeddings using the MOSAIKS method (Rolf et al., 2021).
Key takeaways. The Embed2Scale data challenge at the CVPR EARTHVISION 2025 demonstrated that NeuCo-Bench efficiently evaluates and ranks the performance of compact embeddings over multiple downstream tasks. The scoring method produced a more balanced and discriminative ranking compared to uniformly weighting the tasks, particularly noticeable in the down-weighting of the two random tasks. Hiding the tasks efficiently prevented overfitting, ensuring fairness between the participants. Both winning solutions were based on FMs, indicating these can indeed provide semantically rich, general embeddings. However, also non-FM based solutions scored high.
@@ 5.3 General Evaluations
We assess the NeuCo-Bench framework through a series of experiments spanning three perspectives. First, we follow the challenge setup by constraining all methods to produce 1,024-dimensional embeddings from the input EO data cubes (see Section˜4). We introduce and test approaches including (i) learned neural compressors optimized for rate–distortion trade‐offs, and (ii) self‐supervised FMs, spanning both uni- and multimodal architectures. Second, we relax this requirement to analyze the impact of embedding dimensionality on downstream performance. By varying the size of CNN- and ViT-based embeddings, we investigate whether larger or smaller representations alter relative performance, motivating the 1,024-dimension default and demonstrating the flexibility of the benchmark for exploring embedding size as an additional axis of evaluation. Lastly, we revisit and explore the assumption of linear probes as decoder models.
Data and assumptions. Each input consists of four seasonal snapshots from Sentinel-1 (radar) and Sentinel-2 (optical L1C and L2A). To meet the embedding‐size constraint, we employ frozen, pretrained backbones followed by downsampling, and adopt simplifying assumptions: (1) although S1 and S2 capture complementary signals, substantial correlation justifies unimodal baselines, reflecting common practice in neural compression and FM development on EO data; (2) We fuse the four seasonal views into a single tensor, either before encoding (pre‐encoding aggregation) or after encoding (post‐encoding aggregation), to produce one 1,024‐dimensional vector per sample.
Method Pretraining Loss Backbone Input Downsampling Original latent dim.
TerraMind (FM) TerraMesh Cross-Entropy Multi-modal ViT-B/16 All (S1,S2) Patch averaging $196\times 768$
MAE ViT (FM) SSL4EO MAE ViT-B/16 S2L1C Patch averaging $196\times 768$
DINO ViT (FM) SSL4EO DINO ViT-B/16 S2L1C Patch averaging $196\times 768$
DINO ResNet (FM) SSL4EO DINO ResNet-50 S2L1C Global avg., pairwise channel mean $2048\times 7\times 7$
Factorized Prior SSL4EO Rate-Distortion CNN S2L1C Adaptive pool ($4\times 4$), pairwise channel mean $128\times 14\times 14$
Averaging Baseline All (S1,S2) bilinear avg., flatten $4\times 4\times 8\times 8$
Figure 4: (Left) Radar plot of per‐task $R^{2}$ performance for a rep. subset of embedding–compression methods. Each axis corresponds to one downstream prediction task; the center denotes $R^{2}=0$, negative values are clipped to zero for clarity. (Right) Summary of method configurations.
@@
Methods and baselines. We include a simple averaging baseline, which applies spatial downsampling, channel‐wise averaging, and flattening to yield a 1,024‐dimensional vector; this serves as a minimal reference point. Next, our neural rate–distortion compressors are implemented via Factorized Prior autoencoders (Ballé et al., 2016) pretrained on SSL4EO S2L1C data. We extract latent bottleneck features before entropy coding (cf. (Torfason et al., 2018)) and subsequently downsample them to conform to the 1,024‐dimensional limit. Finally, we evaluate several FMs based on ResNet (He et al., 2015) and ViT (Dosovitskiy et al., 2020) backbones pretrained on SSL4EO data cubes with masked‐autoencoding (MAE (He et al., 2021)) or contrastive (DINO (Caron et al., 2021)) objectives, as well as multimodal architectures like TerraMind (Jakubik et al., 2025), which are pretrained jointly on radar and optical inputs.
@@
Results and discussion. Figure˜4 presents the linear‐probing performance in terms of $R^{2}$ for all downstream tasks assembled and a representative subset of compression methods. Downsampled embeddings from our neural rate–distortion compressors outperform the simple averaging baseline but remain below $R^{2}=0.5$. This observation highlights the characteristics of our setup with high compression rate $\sim$7,000 while testing against linear‐probe. FM embeddings reveal a task‐dependent trend: FMs in general, with contrastive (DINO) and multimodal models (TerraMind) in particular, achieve high $R^{2}$ on semantic tasks where multi-pixel context is relevant (e.g., land‐cover proportion). However, certain FMs struggle on geophysical predictions of quantities resolved at the sub-pixel level (e.g., biomass estimation), whereas multimodal and MAE FMs strike a better balance across tasks.
@@
Temporal aggregation. Across all methods, post‐encoding aggregation consistently outperforms pre‐encoding aggregation. Accordingly, the results in Figure˜4 employ post‐encoding aggregation exclusively, despite incurring a $\times 4$ increase in embedding‐generation runtime. The performance gains are modest for static‐feature tasks (e.g., land‐cover), yet substantial for temporally sensitive tasks (e.g., cloud‐fraction estimation), underscoring the importance of preserving per‐snapshot details.
Embedding size. In Fig.˜5(a) we numerically study how the embedding size impacts performance testing the models introduced above.
@@
  • CNN backbones. Performance peaks for embedding sizes $128\lesssim N\lesssim 1024$, with accuracy dropping outside. Larger embeddings add computational cost without gain in performance.
@@
  • Trade-offs. While larger embeddings increase the number $N=|A|$ of (linear) probe parameters $A$ (cf. Fig.˜1), smaller $N$ often fail to retain task-relevant semantics.
Our experiments support an embedding size of $N=1024$ as a balanced default across downstream tasks, cf. Fig.˜2. The NeuCo-Bench framework is flexible to explore size–utility trade-offs, analogous to rate–distortion analysis in neural compression. Section˜B.3.3 provides additional and detailed per-task performance plots.
((a))
((b))
Figure 5: (a) Average downstream $R^{2}$ as a function of embedding size $\log N$. Largest size is the full channel count (CNN) or native patch-token dimension (ViT). (b) Average downstream $R^{2}$ comparing linear probing against one-layer and two-layer MLP probes on 1024-dimensional embeddings.
Linear probing assumption. We evaluate embedding quality using linear probing, a widely adopted practice in representation learning (Xu and Tewari, 2021) to focus on embeddings without fine-tuning encoder backbones. While non-linear probing (e.g., small MLP heads) can in principle capture richer structures, it risks compensating for poor embedding quality Plachouras et al. (2025). Our experiments in Fig.˜5(b) and Section˜B.3.3 demonstrate: Replacing linear probing by small, non-linear decoders yields only marginal gains for top-performing embeddings, while providing larger improvements for weaker ones. Further, non-linear probing substantially increases computational cost. Thus, linear probing remains an (energy-)efficient and reliable measure of how much semantically relevant information is directly accessible from an embedding space. Linear probing efficiency enabled NeuCo-Bench at the 2025 CVPR EARTHVISION data challenge to run over 400 submissions within minutes on commodity hardware.
@@ 6 Future Work
Reproducibility. Data contributions to NeuCo-Bench demands a permissive CC-BY 4.0 license. Our 2025 CVPR EARTHVISION data challenge required the winners to release their solution under Apache 2.0 license [^3] https://creativecommons.org/licenses/by/4.0 and https://www.apache.org/licenses/LICENSE-2.0.html. The same holds for any future extension of NeuCo-Bench where Section˜B.5 provides further details. While Section˜A.1 has background on the theoretical basis of our evaluation metric with references to code, Sections˜A.2 and B.4 share specifics on the hardware, the software environment, and hyperparameter settings NeuCo-Bench runs on. Section˜A.3 provides some additional elements of running the 2025 CVPR EARTHVISION data challenge. Sections˜B.1 and B.2 gathers general facts on the benchmark data and code framework, respectively. Section˜B.3.2 is dedicated to additional model performance metrics for downstream tasks provided by NeuCo-Bench.
@@
Choice of tasks. The discriminative power of results depends on downstream tasks. We curated diverse, image-level tasks focused on global semantic content, together with a dynamic ranking scheme. Future extensions will include spatially structured tasks such as pixel-level segmentation or time-sensitive predictions, which may require less aggressive compression ratios than the current value of $\sim$7,000. Probing strategies may also evolve as tasks grow in complexity.
@@
Building a NeuCo-Bench community. With NeuCo-Bench, we provide a seed to grow an eco-system centered around benchmarking highly-compressed embeddings on a set of standardized, community-contributed downstream tasks. To avoid contributing to an ever-growing number of benchmark datasets, we intend to harmonize with existing ones, such as GEO-Bench and PANGAEA. NeuCo-Bench’s mission is to provide a standardized framework of benchmarking embeddings. We warmly welcome contributions from all research areas involved in neural compression. Future extensions of the benchmark framework include pixel-wise and temporal downstream tasks.
Societal impact & information privacy. NeuCo-Bench fosters open science for neural compression. SSL4EO-S12 derives from the Copernicus Programme under a free and open data policy. For label sources, please check details in Section˜B.5. The provided dataset is void of detailed time and geolocation information, sparsely spread in time with a single timestamp per quarter, and limited to 10m pixel resolution, limiting any surveillance applications. For benchmarks with strong compression ratio, NeuCo-Bench may act as a privacy conserving approach where pixel-level reconstruction becomes impossible. NeuCo-Bench does not touch on any concerns related to human rights. We include the Linux Foundations Code of Conduct for future contributions to the code released with this paper.

### v2

@@ 5 Benchmark Evaluation
We evaluate the benchmark in two parts. Firstly, Section 5.1 details the real-world evaluation of the benchmark in the 2025 CVPR EarthVision data challenge, as well as discussion of outcomes and learnings in Section 5.2. Secondly, Section 5.3 presents baseline evaluations exploring our set of downstream tasks.
@@ 5.1 Data Challenge Validation
To validate NeuCo-Bench under realistic conditions, we utilized it in the 2025 CVPR EarthVision data challenge. Participants were tasked with compressing multi‐modal, multi‐temporal EO imagery, cf. Section˜4, into 1,024-dimensional embeddings, corresponding to a compression ratio of approximately 7,000. Crucially, participants did not know which or the number of downstream tasks their embeddings would be evaluated on; this hidden‐task design discourages overfitting and encourages the development of general‐purpose EO representations. Participants were ranked, in accordance with Section˜3, across two sets of downstream tasks. One modification was made to the dataset compared to the dataset described in Section˜4; the clouds task targets were mainly replaced by zeros, causing heavy skewness in the labels and basically random connection between the imagery and labels.
Phases. In a three‐week development phase, teams developed embedding methods using the publicly available SSL4EO-S12 v1.1 dataset. A partial release of 5 tasks, each with a subset of its samples but no information of the task type, allowed participants to receive initial feedback for development. Submissions returned only the mean $Q$ value to prevent leakage of task‐specific performance and information.
@@
Platform and infrastructure. The NeuCo-Bench framework was modified to utilize Eval.AI [78] to collect submissions. The benchmark ran on a separate 8-vCPU server, retrieved new submissions via API, executed the evaluation, and pushed results to a custom leader board hosted on GitHub, displaying the dynamic ranking described in Section˜3, as well as back to Eval.AI. Additional details are available in the supplementary material.
@@ 5.2 Data Challenge Results
Participation and ranking. Twenty‐three teams submitted to the development phase; sixteen went on to the final evaluation, nine of which shared their submissions publicly. The quality scores $Q_{t}^{(p)}$, shown in Fig.˜4 for the evaluation phase, varied widely from 0–5 on some tasks to 5–40 on others. The evaluation method, cf. Eq.˜2 , efficiently scaled task importance ensuring that the tasks impacted the leader board relative to their differentiating effect across the teams. Notably, the weighting reduced the impact of the tasks with random labels. Further, the dynamic ranking caused a swap between the original first and second place methods due to a third team, highlighting the impact of adaptive weighting. Detailed specifications of the challenge evaluations is given in the supplementary material.
Figure 4: Quality score $Q_{t}^{(p)}$ of the participants of the data challenge evaluation phase with corresponding task weight used for ranking. Teams are ordered by final leader board rank (Team 1 is the winner). Team 10 is a simple averaging baseline (Section˜5.3) and Team 11 is a normally distributed random embeddings baseline. We observe that landcover tasks are the most discriminative, the random control task is the most “difficult”(by design), and estimating biomass is the most challenging real-world downstream task.
Top methods. The team that achieved the best overall rank and the team with the highest average $Q$-Score both built their embeddings by ensembling multiple FM representations, although with different approaches; One pre-training backbones, and the other training a bottleneck based on frozen FM backbones. The fourth-place team took a different path, forgoing any pre-training and instead generating embeddings using the MOSAIKS method [54].
Key takeaways. The EV challenge demonstrated that NeuCo-Bench efficiently evaluates and ranks the performance of compact embeddings over multiple downstream tasks. The scoring method produced a more balanced and discriminative ranking compared to uniformly weighting the tasks, particularly noticeable in the down-weighting of the two random tasks. Hiding the tasks efficiently prevented overfitting, ensuring fairness between the participants. Both winning solutions were based on FMs, indicating these can indeed provide semantically rich, general embeddings. However, also non-FM based solutions scored high.
@@ 5.3 General Evaluations
We assess the NeuCo-Bench framework through a series of experiments, including embeddings from self-supervised FMs and representations from learned neural compressors. To meet the compression requirements, we apply spatial and temporal aggregation techniques to obtain compact embeddings.
Our analyses span three perspectives: (i) we follow the challenge setup and constrain all methods to 1,024-dimensional embeddings; (ii) we relax the embedding dimension constraint to study how embedding dimensionality affects downstream performance, motivating the challenge default; and (iii) we revisit the assumption of linear probes replacing them by more complex decoder models.
Embedding Aggregation Methods. Each input consists of four seasonal snapshots from Sentinel-1 (radar) and Sentinel-2 (optical), as described in Section˜4. We benchmark image-based encoders by applying temporal averaging across the four timesteps, either before encoding (pre-encoding aggregation) or after encoding (post-encoding aggregation). Convolutional encoder outputs are aggregated through spatial averaging in embedding space, with additional pairwise channel means applied when further dimensionality reduction is needed. For ViT encoders, we either average the spatial patch tokens to obtain a single embedding or use the CLS-token. Aggregated embeddings smaller than the target size (e.g., 1024) are zero-padded. To study the role of input modalities, we evaluate unimodal encoders that use only S2L1C bands and multimodal encoders that process the full SSL4EO-S12 data cube.
Table 2: Full per‐task $R^{2}$ scores for tested embedding methods including average performance (Avg.) across all tasks. For ViT’s, experiments utilizing the CLS token for prediction is indicated by (CLS), otherwise the average patch token is used. Methods are sorted in ascending order by Avg. For each task, the best-performing method is highlighted in bold, and the second‐best method is underlined. Abbreviations: Landcover (LC), Heatisland (HI), Agriculture (Agri), Average (Avg.) across all downstream tasks
Method Biomass Mean Biomass Std Crops Clouds LC Agri LC Forest HI Mean HI Std Avg.
Averaging Baseline -0.552 -0.426 -0.061 -2.293 0.126 0.216 -0.962 -1.157 -0.514
Factorized Prior ($\lambda$ = 0.5) -0.036 -0.071 0.325 0.200 0.478 0.343 0.033 -0.628 0.080
Factorized Prior ($\lambda$ = 0.1) 0.129 0.078 0.357 0.266 0.464 0.333 0.219 -0.270 0.197
Factorized Prior ($\lambda$ = 0.025) 0.195 0.140 0.338 0.288 0.449 0.256 0.315 -0.113 0.233
S12-DINO ResNet50 [71] 0.117 0.088 0.826 0.147 0.879 0.865 0.483 -0.221 0.286
S12-MoCo ResNet50 [71] -0.048 -0.030 0.780 0.216 0.864 0.844 0.345 -0.334 0.286
Satlas MS ResNet50 [7], 0.310 0.248 0.490 0.340 0.658 0.641 0.485 0.092 0.408
Clay-V1 ViT-B (CLS) [76] 0.199 0.169 0.748 0.482 0.811 0.803 0.430 -0.013 0.454
S12-DINO ViT-S [71] 0.184 0.183 0.842 0.484 0.863 0.851 0.461 -0.109 0.470
S12-MoCo ViT-S [71] 0.338 0.259 0.751 0.409 0.825 0.814 0.506 0.133 0.504
DOFA-ViT-L [76] 0.373 0.269 0.587 0.518 0.777 0.773 0.560 0.207 0.508
S12-MAE ViT-S [71] 0.380 0.280 0.627 0.659 0.792 0.797 0.585 0.181 0.537
Prithvi-EO-V2-300M (CLS) [62] 0.426 0.329 0.655 0.531 0.764 0.782 0.603 0.216 0.538
Prithvi-EO-V2-300M [62] 0.424 0.324 0.671 0.588 0.781 0.792 0.608 0.203 0.549
Clay-V1 ViT-B [76] 0.466 0.332 0.759 0.660 0.825 0.825 0.606 0.229 0.588
TerraMind-V1-B [35] 0.528 0.390 0.879 0.731 0.918 0.908 0.691 0.226 0.659
@@
Encoder Baselines. A simple averaging baseline, which applies spatial pooling, channel‐wise averaging, and flattening serves as a minimal reference point, see the supplementary for details. Neural rate–distortion compressors are implemented via Factorized Prior autoencoders [5] pretrained on SSL4EO S2L1C data. We extract latent bottleneck features before entropy coding (cf. [64]) and aggregate as described above. We benchmark a range of publicly available EO FMs, including unimodal ResNet [31] and ViT [17] backbones, and multimodal models like TerraMind [35] and DOFA [76].
@@
Results and discussion. Table˜2 presents the linear‐probing performance in terms of $R^{2}$ for all downstream tasks. Embeddings from neural rate–distortion compressors outperform the simple averaging baseline but generally remain below $R^{2}=0.5$. This reflects the characteristics of our setup which combines high compression rate $\sim$7,000 with linear probing. FM embeddings show a task‐dependent trend: FMs in general, with contrastive (DINO, MoCo) and multimodal models (TerraMind) in particular, achieve high $R^{2}$ on semantic tasks where multi-pixel context is relevant (e.g., land‐cover proportion). However, certain FMs struggle on geophysical predictions of quantities resolved at the sub-pixel level (e.g., biomass estimation).
@@
Temporal aggregation. Across all methods, post‐encoding aggregation consistently outperforms pre‐encoding aggregation (see the supplementary). Accordingly, the results in Table˜2 all use temporal post‐encoding aggregation. The gains are modest for static‐feature tasks (e.g., land‐cover), yet substantial for temporally sensitive tasks (e.g., cloud‐fraction estimation), underscoring the importance of preserving per‐snapshot details.
Embedding size. In Fig.˜5 we study numerically how embedding size impacts performance on a subset of models, with per-task results in the supplementary material. Given our setup, we observe:
@@
  • CNN backbones. Performance peaks for $128\lesssim N\lesssim 1024$ and drops outside this range. Larger embeddings add computational cost without significant performance gains.
@@
  • Trade-offs. While larger embeddings increase the number $N=|A|$ of (linear) probe parameters $A$ (cf. Fig.˜2), smaller $N$ often fail to retain task-relevant semantics.
Figure 5: Average task $R^{2}$ vs. embedding size $\log N$. Largest size is the full channel count (CNN) or patch-token dim. (ViT).
Our experiments support an embedding size of $N=1024$ as a practical default across tasks, cf. Fig.˜3. The NeuCo-Bench framework is flexible to explore size–utility trade-offs, analogous to rate–distortion analysis in neural compression.
Linear probing assumption. We evaluate embedding quality using linear probing, a standard practice in representation learning [77], to assess embeddings without fine-tuning encoder backbones. While non-linear probing (e.g., small MLP heads) can capture richer structures, they can compensate for poor embedding quality [49].
Figure 6: Average $R^{2}$ comparing linear probing against one-layer and two-layer MLP probes on 1024-dimensional embeddings.
Our experiments in Fig.˜6 and with additional details in the supplementary material show that replacing linear probes by small, non-linear decoders yields only marginal gains for top-performing embeddings, while providing larger improvements for weaker ones, at substantially higher computational cost. Linear probing remains an efficient and reliable measure of how much semantically relevant information is directly accessible from an embedding space. Its efficiency enabled NeuCo-Bench at the EV challenge to run over 400 submissions within minutes on commodity hardware.
@@ 6 Future Work
Reproducibility. Data contributions to NeuCo-Bench demands a permissive CC-BY 4.0 license. Our data challenge required the winners to release their solution under Apache 2.0 license [^3] https://creativecommons.org/licenses/by/4.0 and https://www.apache.org/licenses/LICENSE-2.0.html. The same holds for any future extension of NeuCo-Bench. Details how to extend the framework, reference code, hardware specifications, software environments and hyperparameters used in the data challenge are provided in the supplementary material.
@@
Choice of tasks. The discriminative power of results depends on downstream tasks. We curated diverse, image-level tasks focused on global semantic content, together with a dynamic ranking scheme. Future extensions will include spatially structured tasks such as pixel-level segmentation or time-sensitive predictions, which require less aggressive compression ratios than the current value of $\sim$7,000. Probing strategies may also evolve as tasks grow in complexity.
@@
Building a NeuCo-Bench community. With NeuCo-Bench, we provide a seed to grow an ecosystem centered around benchmarking compact embeddings on a set of standardized, community-contributed downstream tasks. To avoid contributing to an ever-growing number of benchmark datasets, we intend to harmonize with existing ones, such as GEO-Bench and PANGAEA. Future extensions of the benchmark framework include pixel-wise and temporal downstream tasks.
Societal impact & information privacy. NeuCo-Bench fosters open science for neural compression. SSL4EO-S12 derives from the Copernicus Programme under a free and open data policy. For label sources, please check details in the supplementary material. The provided dataset is void of detailed time and geolocation information, sparsely spread in time with a single timestamp per quarter, and limited to 10m pixel resolution, limiting any surveillance applications. For benchmarks with strong compression ratio, NeuCo-Bench may act as a privacy conserving approach where pixel-level reconstruction becomes impossible. NeuCo-Bench does not touch on any concerns related to human rights. We include the Linux Foundations Code of Conduct for future contributions to the code released with this paper.